OpenStudy (anonymous):

find two linearly independent solutions to the differential equation (3x-1)^2y"+(9x-3)y'-9y=0

OpenStudy (anonymous):

Not sure that this is what is required. From Mathematica v9:

OpenStudy (michele_laino):

If we change variable, namely: \[\large 3x - 1 = s\] where s is a new variable of differentiation, we can rewrite your differential equation as below: \[\large {s^2}{y_{ss}} + s{y_s} - y = 0\] where \[\large {y_s} = \frac{{dy}}{{ds}},\;{y_{ss}} = \frac{{{d^2}y}}{{d{s^2}}}\] Now, that last form of your equation is a homogeneous Euler equation. In order to find its solutions, we try this substitution: \[\large y = k{s^a}\] After that substitution, we get: \[\large a = 1,\quad a = - 1\] (please check those values) so two linearly independent solutions, are: \[\large y = s,\quad y = \frac{1}{s}\] or using the definition of s: \[\large {y_1} = 3x - 1,\quad {y_2} = \frac{1}{{3x - 1}}\]